1. The data shown below are the price (in £s) of a ticket for a sample of 25 music concerts. Find the mode, mean and the median for this data set.
  2. 27 22 63 65 20 45 47 33 96 55 12 45 54 48 73
  3. 8 89 58 19 43 15 17 94 37 21
1
Expert's answer
2020-12-03T12:39:38-0500

Mode is the value with highest frequency which is 45

Mean = ∑X÷N = 1106÷25 = 44.24

Median is the central value in a series (8, 12, 15, 17, 19, 20, 21, 22, 27, 33, 37, 43, 45, 45, 47, 48, 54, 55, 58, 63, 65, 73, 89, 94, 96) which is the 13th value in the series.

Therefore the median is 45